# High Frequency Quoting Under Liquidity Constraints

### v1

@@ 4.3 Execution using limit orders
The interval $[t_{1},t_{2}]$ is referred to as the critical interval.
Notation Convention:
@@
The time difference between the execution of the quoting leg and the reference leg is called the critical interval. During this interval, the top-of-book price on the reference leg may change. This affects the realized spread $\tilde{S}^{(r)}(t_{2})$, causing deviation from the trader’s mandated spread $S(t_{0})$. We define the slippage incurred as:
@@
Top of Book at $t_{0}$
Table 1: Top 5 levels of the LOB at time $t_{0}$ for NIFTY February and March Futures
NIFTY February 2022
@@
NIFTY March 2022
@@
From this, the two implied spreads are:
@@
In this setup, the trader designates the current month contract as the reference leg. A buy limit order is placed on the next month contract at time $t_{0}$, quoting the best available ask price of $p^{a}_{n}(t_{0})=17514.2$. This order is executed at time $t_{1}$.
Following this execution, at time $t_{2}$, the trader fires an aggressive limit order to sell the current month contract. During the critical interval $[t_{1},t_{2}]$, the top bid on the current month leg improves from 17455 to 17458. Consequently, the aggressive order is executed at the updated price $p^{b}_{c}(t_{2})=17458$.
The spread realized under this execution path, denoted $\tilde{S}^{(c)}(t_{2})$, is calculated as the difference between the bid on the next month leg (observed at $t_{1}$) and the updated bid on the reference leg (current month) at $t_{2}$. This spread is 56.2 INR, leading to a slippage of 2.3 INR compared to the mandated spread of 58.5 INR.
@@
Table 2: LOB snapshots when current month is the reference leg
NIFTY March 2022 ($t_{1}$)
@@
In this alternative configuration, the trader chooses the next month contract as the reference leg. A buy limit order is placed on the current month contract at time $t_{0}$, quoting the best available ask price of $p^{a}_{c}(t_{0})=17458.55$. This order is filled at time $t_{1}$.
At $t_{2}$, the trader fires an aggressive limit order on the next month contract to sell at market. During the critical interval, the top ask price on the next month contract shifts slightly from 17516.5 to 17516.55, reflecting minor upward movement. The order is thus executed at the updated price $p^{a}_{n}(t_{2})=17516.55$.
The realized spread $\tilde{S}^{(n)}(t_{2})$, computed as the difference between the ask on the reference leg (next month) at $t_{2}$ and the ask on the quoting leg (current month) at $t_{1}$, equals 58.0 INR. This results in a slippage of only 0.5 INR from the mandated spread.
@@
Table 3: LOB snapshots when next month is the reference leg
NIFTY February 2022 ($t_{1}$)
@@
Table 4: Reference Leg Pricing Path
Ref. Leg Quoted Price Exec. Price Ref Price ($t_{0}$) Ref Exec ($t_{2}$)
Current Month ($r=c$) 17514.2 17514.2 17455 17458
Next Month ($r=n$) 17458.55 17458.55 17516.5 17516.55
Table 5: Realized Spread and Slippage
Ref. Leg Realized Spread Slippage (INR)
Current Month ($r=c$) 56.2 2.3
Next Month ($r=n$) 58.0 0.5
@@
This highlights the core insight: short-term stability of the reference leg is more important than instantaneous spread width when minimizing execution risk.
@@ 5 Problem Formulation: Optimal Reference Leg Selection
The preceding empirical comparison illustrates that the choice of reference leg in a calendar spread quoting strategy materially affects the realized spread and slippage. Specifically, even when the initial observed spread favors one contract, short-term price stability during the critical interval can dominate realized execution outcomes. Motivated by these observations, we now formalize the reference leg selection problem as a dynamic decision-making process. At each quoting decision point, the trader must select the contract to treat as the reference leg based on observable market conditions, with the objective of minimizing expected slippage relative to a mandated spread constraint. Unlike classical optimization problems, this selection is guided by statistical methods and empirical characteristics of the limit order book rather than by solving an explicit minimization problem. This formulation captures the execution asymmetry inherent in quoting strategies and provides the foundation for systematic reference leg selection based on real-time market data.
@@
  • Filters $\mathcal{G}^{\mathcal{T}}(w)$ that act on the stream of trades and quotes observed strictly prior to the start of the interval.
  • Filters $\mathcal{G}^{\mathcal{B}}(w)$ that operate on the order book snapshot observed at the start of the interval.
@@
Each filter is paired with a quoting function that maps the filtered representation to a real-valued score for each candidate leg. The tick-based quoting function $\mathcal{Q}^{\mathcal{T}}$ operates on the output of $\mathcal{G}^{\mathcal{T}}(w)$, while the book-based quoting function $\mathcal{Q}^{\mathcal{B}}$ operates on the output of $\mathcal{G}^{\mathcal{B}}(w)$.
@@
We denote by $\chi^{m}(w)\in\{0,1\}$ the market-optimal reference leg decision for interval $w$, defined as the choice that yields the lowest realized slippage in hindsight. This benchmark provides a reference against which to evaluate the quality of decisions made by each quoting rule.
Problem 1.2 (Evaluation of Reference Leg Selection Rules). Given a collection of evaluation intervals $\mathcal{W}$, and binary-valued quoting decisions $\chi^{\mathcal{T}}(w)$, $\chi^{\mathcal{B}}(w)$, and the market-optimal reference decision $\chi^{m}(w)\in\{0,1\}$, define the agreement score: $\mathcal{A}(\chi,\chi^{m}):=\frac{1}{|\mathcal{W}|}\sum_{w\in\mathcal{W}}\mathbb{I}\left\{\chi(w)=\chi^{m}(w)\right\}$ The objective is to compute and compare the agreement scores $\mathcal{A}(\chi^{\mathcal{T}},\chi^{m})$ and $\mathcal{A}(\chi^{\mathcal{B}},\chi^{m})$.
@@ 6 Methodology
The decision rules introduced in Problem 1.2 are evaluated against an oracle benchmark that reflects the most favorable realized execution observed in hindsight. In the context of machine learning, an oracle typically refers to an idealized entity or function that provides the correct label or decision for each input. Here, the oracle decision $\chi^{m}(w)$ identifies the reference leg—either current or next month—that would have minimized realized slippage within a given evaluation window $w$. It serves as a ground truth label for scoring the performance of quoting strategies based on observable signals. The methodology outlined in this section implements the decision rules $\chi^{\mathcal{T}}(w)$ and $\chi^{\mathcal{B}}(w)$, instantiates the oracle decision $\chi^{m}(w)$ from tick data, and computes the agreement score defined in Problem 1.2 as the primary evaluation metric.
@@ 6.1 Execution Design and Quoting Framework
Let $F_{c}$ be the futures contract expiring in the current month (with price $p_{c}(t)$ at time $t$), and let $F_{n}$ be the contract expiring in the next month (with price $p_{n}(t)$). Let $q$ denote the quantity to roll over, and $S$ denote the maximum acceptable rollover cost (price difference per unit). Executing a rollover trade then entails selling $q$ units of $F_{c}$ and buying $q$ units of $F_{n}$, under the constraint $p_{n}(t)-p_{c}(t)\leq S$, ensuring total cost remains within the threshold $S$.
In an efficient, arbitrage-free market with risk-free rate $R$, the prices satisfy the theoretical cost-of-carry relation: $p_{n}(t)=e^{Rt}p_{c}(t)$. This leads to the theoretical spread:
$S_{m}(t)=p_{c}(t)(e^{Rt}-1),$
which may deviate in practice due to mispricing or market frictions. These frictions motivate quoting strategies that dynamically adjust order prices to maintain execution costs below $S$.
@@
The price $p_{n}^{r}(t)$ is derived similarly from the ask side of $F_{n}$’s LOB:
@@
with quoting price on $F_{c}$ set to:
$p_{c}^{l}(t)=p_{n}^{r}(t)-S.$
The realized spread $\tilde{S}(t)$ is computed from the resulting trade prices after quoting and aggressive leg execution. A key design decision is determining which contract to use as the reference leg, since this impacts execution cost and slippage.
6.2 Tick-Based and Book-Based Reference Leg Selection
@@
Tick-based quoting rule.
@@
Book-based quoting rule.
Similarly, for each $w\in\mathcal{W}$, we define a book-based quoting rule that uses the order book snapshot observed at time $w_{1}$ to evaluate expected slippage:
with associated decision function:
These rules implement quoting strategies presented in 5 based on real-time market observables, where tick-based filters aggregate historical event data, and book-based filters extract instantaneous liquidity conditions. The decision outcome in each case is a binary indicator for whether to quote on the next-month leg.
7 Benchmarking Framework for Quoting Strategies
Let =$\mathcal{E}=\{T,C,\text{PDM}\}$ denote the set of event types observed in tick data, where $T$ represents trades, $C$ represents cancellations, and PDM denotes price downward modifications.
[…]
Figure 1: Illustration of benchmark construction using passive-aggressive trade pairs within interval $w=[w_{1},w_{2})$. The first-leg execution determines the oracle decision $\chi^{m}(w)$.
Remark 4.
(Trade-Only Basis for Benchmark Construction)
Only Trade events are used to construct the benchmark decision $\chi^{m}(w)$. Events of type Cancellation or Price Downward Modification are excluded from the filtered set $\mathcal{G}^{\mathcal{T}}(w)$, ensuring that benchmarking reflects executed spread trades alone.
$\tilde{S}_{m}(t)=p_{n}^{b,T}(t_{2})-p_{c}^{a,T}(t_{1})$
, where $T$ represents event-type Trade. The oracle decision is:
$\chi^{m}(w):=\begin{cases}1&\text{if }F_{c}\text{ was traded first (ref leg is}F_{n})\\ 0&\text{if }F_{n}\text{ was traded first (ref leg is }F_{c})\end{cases}$
Remark 5.
(On Quantity Normalization)
In plain terms, the oracle label $\chi^{m}(w)$ is determined by finding actual calendar spread trades in the tick data. These trades are identified as passive-aggressive trade pairs in opposite contracts, occurring within a short critical interval. We retain only those trades that match rollover logic—buying $F_{n}$, selling $F_{c}$—and compute the realized spread for each. The lowest spread in the window determines the oracle decision based on which leg executed first. The complete filtered set of such trades is denoted $\mathcal{G}^{\mathcal{T}}(w)$, and all benchmarking is grounded in this empirically observed structure.
8 Data
This study uses tick-by-tick (TBT) market data from the National Stock Exchange (NSE) of India, collected on February 7, 2022, for NIFTY futures contracts.  [^3] All exchange specifications and market microstructure rules referenced in this study reflect those in effect as of March 2022, corresponding to the period from which the data was collected. The dataset includes two instruments: the current-month contract expiring on February 24, 2022 ($F_{c}$) and the next-month contract expiring on March 31, 2022 ($F_{n}$). Both contracts exhibit high liquidity and dense tick arrival activity, enabling fine-grained execution modeling.
Each tick is one of four event types: NEW (order addition), MODIFY (order revision), CANCEL (order withdrawal), and TRADE (execution). These types are used to extract features relevant to quoting behavior and limit order book dynamics. The minimum tick size for NIFTY futures is INR 0.05, as mandated by NSE.
The data is segmented into a collection of evaluation windows $\mathcal{W}=\{w=[w_{1},w_{2})\}$, where each window spans a short time interval (e.g., 100 milliseconds) aligned with the quoting decision frequency. For each $w\in\mathcal{W}$, we apply two distinct classes of filters:
  • The tick-based filter $\mathcal{G}^{\mathcal{T}}(w)$ operates on the stream of trades and quotes observed strictly prior to the quoting time $w_{1}$.
  • The book-based filter $\mathcal{G}^{\mathcal{B}}(w)$ extracts features from the limit order book snapshot observed exactly at time $w_{1}$.
Tick-derived features.
For each evaluation window $w\in\mathcal{W}$, the tick-based filter $\mathcal{G}^{\mathcal{T}}(w)$ encodes summary statistics of recent trade and quote activity observed strictly prior to the quoting time $w_{1}$. These features include:
  • Count of executed trades (TRADE events).
  • Count of order cancellations (CANCEL events).
  • Count of price-worsening quote updates (MODIFY events where the best price deteriorates).
  • Total number of tick events, inclusive of all above types.
Book-derived features.
For each interval $w\in\mathcal{W}$, the book-based filter $\mathcal{G}^{\mathcal{B}}(w)$ extracts directional liquidity indicators from the limit order book snapshot observed at time $w_{1}$. These features are computed separately for each contract leg and for each relevant side of the book (bid or ask), based on whether the leg is expected to be lifted or hit under the quoting convention.
Specifically, we extract:
  • Price at each level $p^{b,i},p^{a,i}$, for $i=1,\dots,\nu$
  • Quantity available at each level $q^{b,i},q^{a,i}$, for $i=1,\dots,\nu$
These level-wise features are later used to construct directional liquidity measures such as the Composite Liquidity Factor (CLF), introduced in a subsequent section, which serve as inputs to the book-based quoting rule $\mathcal{Q}^{\mathcal{B}}$.
Data scale.
Table 6 reports the total number of ticks recorded for each contract on February 7, 2022. The substantial volume supports the evaluation of quoting strategies at a high temporal resolution.
Contract Date Number of ticks/day
Futures (G22) 07/02/2022 49,828,303
Futures (H22) 07/02/2022 16,349,197
Futures (H22) 10/03/2022 42,000,613
Futures (J22) 10/03/2022 9,562,649
Futures (H22) 24/03/2022 43,306,045
Futures (J22) 24/03/2022 18,238,164
Table 6: Data Statistics (The code in the parenthesis denotes expiry in the format: ⟨month_code⟩⟨year⟩)
A sample from the processed tick data is provided in the appendix for illustration.
9 Forecasting Execution Risk with Multivariate Hawkes Processes

### v2

@@ 1.4 Impact of Reference Leg Selection on Execution Slippage
The interval $t_{1},t_{2}]$ is referred to as the critical interval. We reiterate here for clarity, the notation Convention:
@@
The time difference between the execution of the quoting leg and the reference leg is called the critical interval. During this interval, the top-of-book price on the reference leg may change. This affects the realized spread $\tilde{S}^{(r)}(t_{2})$, causing deviation from the trader’s target spread $S(t_{0})$. We define the slippage incurred as:
@@
Top of the Book at $t_{0}$
(a) NIFTY February 2022 (Prices in INR; sizes in contracts)
@@
(b) NIFTY March 2022 (Prices in INR; sizes in contracts)
@@
Table 1: Top 5 LOB levels at $t_{0}$ for NIFTY futures
From the limit order books in table 1, at time $t_{0}$, the two implied spreads are:
@@
In this setup, the trader designates the current-month contract as the reference leg. A buy limit order is placed on the next-month contract at time $t_{0}$, quoting the best available ask price of $p^{a}_{n}(t_{0})=17514.2$. This order is executed at time $t_{1}$.
Following this execution, at time $t_{2}$, the trader fires a marketable limit order to sell the current-month contract. During the critical interval $t_{1},t_{2}]$, the top bid on the current month leg improves from 17455 to 17458. Consequently, the aggressive order is executed at the updated price $p^{b}_{c}(t_{2})=17458$.
The spread realized under this execution path, denoted $\tilde{S}^{(c)}(t_{2})$, is calculated as the difference between the bid on the next month leg (observed at $t_{1}$) and the updated bid on the reference leg (current month) at $t_{2}$. This spread is —56.2, leading to a slippage of —2.3 compared to the target spread of —58.5.
@@
(a) NIFTY March 2022 ($t_{1}$)
@@
Table 2: LOB snapshots when current month is the reference leg
@@
In this alternative configuration, the trader chooses the next-month contract as the reference leg. A buy limit order is placed on the current-month contract at time $t_{0}$, quoting the best available ask price of $p^{a}_{c}(t_{0})=17458.55$. This order is filled at time $t_{1}$.
At $t_{2}$, the trader fires a marketable limit order on the next-month contract to sell at market. During the critical interval, the top ask price on the next-month contract shifts slightly from 17516.5 to 17516.55, reflecting minor upward movement. The order is thus executed at the updated price $p^{a}_{n}(t_{2})=17516.55$.
The realized spread $\tilde{S}^{(n)}(t_{2})$, computed as the difference between the ask on the reference leg (next month) at $t_{2}$ and the ask on the quoting leg (current month) at $t_{1}$, equals —58.0. This results in a slippage of only —0.5 from the target spread.
@@
(a) NIFTY February 2022 ($t_{1}$)
@@
Table 3: LOB snapshots when next month is the reference leg
@@
Ref. Leg Ref Price ($t_{0}$) Ref Exec ($t_{2}$) Realized Spread Slippage (INR)
Current Month ($r=c$) 17455 17458 56.2 2.3
Next Month ($r=n$) 17516.5 17516.55 58.0 0.5
Table 4: Reference leg outcomes: reference price at $t_{0}$ and execution at $t_{2}$ with realized spread and slippage (in INR) for the current month $(r=c)$ and next month $(r=n).$
@@
This highlights the core insight: short-term stability of the reference leg is more important than instantaneous spread width when minimizing execution risk. In the sections that follow, we derive leg-specific stability estimates from LOB state (CLF) and from event history via a Hawkes arrival ratio, and use them to evaluate reference-leg choices.
@@ 2 Problem Formulation: Optimal Reference Leg Selection
The preceding empirical comparison illustrated that the choice of reference leg in a calendar spread quoting strategy materially affects the realized spread and slippage. Specifically, even when the initial observed spread favors one contract, short-term price stability during the critical interval can dominate realized execution outcomes. Motivated by these observations, we now formalize the reference leg selection problem as a dynamic decision-making process. At each quoting decision point, the trader must select the contract to treat as the reference leg based on observable market conditions, to minimize expected slippage relative to a target spread constraint. Unlike classical optimization problems, this selection is guided by statistical methods and empirical characteristics of the limit order book rather than by solving an explicit minimization problem. This formulation captures the execution asymmetry inherent in quoting strategies and provides the foundation for systematic reference leg selection based on real-time market data.
@@
  1. 1.
    Filters $\mathcal{G}^{\mathcal{T}}(w)$ that act on the stream of trades and quotes observed strictly before the start of the interval.
  2. 2.
    Filters $\mathcal{G}^{\mathcal{B}}(w)$ that operate on the LOB snapshot observed at the start of the interval.
@@
Each filter is paired with a quoting function that maps the filtered representation to a real-valued score for each candidate leg. The tick-based quoting function $\mathcal{Q}^{\mathcal{T}}$ operates on the output of $\mathcal{G}^{\mathcal{T}}(w)$, while the LOB-based quoting function $\mathcal{Q}^{\mathcal{B}}$ operates on the output of $\mathcal{G}^{\mathcal{B}}(w)$.
@@
We denote by $\chi^{m}(w)\in\{0,1\}$ the market-optimal reference leg decision for interval $w$, defined as the choice that yields the lowest realized slippage in hindsight. This oracle benchmark provides a reference against which to evaluate the quality of decisions made by each quoting rule.
Remark.
We report realized spread slippage because the remaining cost components (exchange fees, rebates, and longer–horizon opportunity costs) are invariant to the reference–leg choice under our setup: a single venue and identical contract multipliers. These terms therefore add an interval–level constant and do not affect the ranking of $\chi^{\mathcal{B}}$ versus $\chi^{\mathcal{H}}$. In market settings with asymmetric fee schedules, cross–venue routing, or role–dependent fills, these terms need not be invariant. Integrating them is orthogonal to our information comparison and is left for future work.
Problem Definition: Evaluation of Reference Leg Selection Rules. Let $\mathcal{W}=\{w\}$ be non–overlapping evaluation intervals $w=[w_{1},w_{2})$. For each $w$, let $\chi^{\mathcal{B}}(w),\chi^{\mathcal{H}}(w)\in\{0,1\}$ denote the binary decisions produced at $w_{1}$ by the CLF–based rule and the Hawkes–based rule, respectively (value 11 selects the next–month leg, value 0 selects the current–month leg). Let $\delta_{S}^{(r)}(w_{2})$ be realized slippage on leg $r\in\{c,n\}$ over $w_{1},w_{2})$. Given the market–optimal oracle decision $\chi^{m}(w)\;:=\;\begin{cases}1,&\text{if }\delta_{S}^{(n)}(w_{2})<\delta_{S}^{(c)}(w_{2}),\\ 0,&\text{if }\delta_{S}^{(c)}(w_{2})\leq\delta_{S}^{(n)}(w_{2}),\end{cases}$ which breaks ties in favor of $c$ deterministically, we define the agreement score: $\mathcal{A}(\chi,\chi^{m}):=\frac{1}{|\mathcal{W}|}\sum_{w\in\mathcal{W}}\mathbb{I}\left\{\chi(w)=\chi^{m}(w)\right\}$ Compute and compare the agreement scores $\mathcal{A}(\chi^{\mathcal{T}},\chi^{m})$ and $\mathcal{A}(\chi^{\mathcal{B}},\chi^{m})$. Note: Decisions $\chi^{\mathcal{B}}(w)$ and $\chi^{\mathcal{H}}(w)$ are formed using only information timestamped strictly before $w_{1}$. The realized slippage on $w_{1},w_{2})$ is used only for evaluation and to define $\chi^{m}(w)$.
@@ 3 Methodology
The decision rules introduced in Section 2 are evaluated against an oracle benchmark. In this paper, the oracle is an infeasible hindsight lower bound on deviation from the target spread: for each interval $w$, the oracle decision $\chi^{m}(w)$ selects the leg (current or next month) that would have minimized realized slippage on $w_{1},w_{2})$. We use this hindsight choice as an evaluation label for scoring data-driven quoting rules formed from observable signals. The methodology implements $\chi^{\mathcal{T}}(w)$ and $\chi^{\mathcal{B}}(w)$, instantiates $\chi^{m}(w)$ from realized outcomes, and computes the agreement score defined in the Problem Definition of Section 2 as the primary evaluation metric. All decisions use only information timestamped strictly before $w_{1}$; outcomes on $w_{1},w_{2})$ are used solely for evaluation.
@@ 3.1 Execution Design and Quoting Framework
This subsection specifies how the executable quoting price is computed after the reference leg is chosen: we restate the notation for completeness, benchmark against the no-arbitrage spread, and then express the executable price as a VWAP on the selected reference leg. We frame the problem as a calendar-spread rollover. A common configuration is buying the spread: selling the current-month contract and buying the next-month contract simultaneously. Execution requires a side-conditional decision—whether the bid in the current-month or the ask in the next-month offers better instantaneous liquidity for the required size. We treat this as a reference-leg selection problem and evaluate the choice using deviation from the target spread as per the oracle benchmark (Section 3.6).
Let $F_{c}$ denote the futures contract expiring in the current-month (priced $p_{c}(t)$ at time $t$) and $F_{n}$ the next-month contract (priced $p_{n}(t)$). Let $q$ be the quantity to trade and let $S$ denote a fixed target-spread threshold (a maximum acceptable cost per unit, distinct from the time-varying spread $S(t)=p_{n}(t)-p_{c}(t)$). A calendar spread trade sells $q$ units of $F_{c}$ and buys $q$ units of $F_{n}$ subject to the constraint $p_{n}(t)-p_{c}(t)\leq S$. We next benchmark this threshold against the theoretical spread implied by no-arbitrage carry. Conditional on the reference-leg decision, the executable quoting price is taken as the volume-weighted average of the quoted prices required to fill $q$ on the chosen leg at time $t$. The next paragraph makes this VWAP construction explicit.
In an efficient, arbitrage-free market with risk-free rate $R$ and identical multipliers on the current- and next-month contracts (ignoring dividends/funding spreads), the cost-of-carry relation implies $p_{n}(t)=e^{R(t_{n}-t_{c})}\,p_{c}(t)$, where $t_{c}$ and $t_{n}$ denote the expiry times of the current- and next-month contracts, respectively, and $t_{n}-t_{c}$ is measured in years consistent with the compounding convention of $R$. This yields the model-implied spread
$S_{m}(t)=p_{c}(t)\,\big(e^{R(t_{n}-t_{c})}-1\big)$
. which may deviate in practice due to mispricing or market frictions. These frictions motivate quoting policies that dynamically adjust order prices to maintain execution costs below $S$.
@@
Conditional on the reference-leg choice, we now compute the executable quoting price on the chosen leg using a VWAP of the LOB levels needed to fill $q$ at time $t_{0}$.
@@
Conditional on selecting the next-month contract as the reference leg, the reference leg is executed via a marketable limit order at time $t$. The executable reference price $p_{n}^{r}(t)$ is taken as the VWAP of the ask-side levels of the limit order book of $F_{n}$ required to fill the target size $q$:
@@
where $\nu$ is the smallest depth index such that $\sum_{i=1}^{\nu}q_{n}^{a,i}\geq q$; the last level is partially consumed if needed. The linked quoting price on the current-month contract is then set to
$p_{c}^{l}(t)=p_{n}^{r}(t)-S,$
so that the quoted spread equals the target $S$ under the convention $S(t)=p_{n}(t)-p_{c}(t)$. Here, the superscripts $r$ and $l$ denote the reference-leg VWAP and the linked quote, respectively. (If the trading direction requires selling $F_{n}$, the construction is symmetric with bid-side levels.)
The realized spread $\tilde{S}(t)$ is computed from the resulting trade prices after quoting and executing the reference leg with a marketable limit order. Consequently, the choice of reference leg affects execution cost and slippage, and will be evaluated against the oracle benchmark in the next subsection.
3.2 Benchmarking Framework for quoting policies
Let $\mathcal{E}=\{T,C,\text{PDM}\}$ denote the set of event types observed in tick data, where $T$ represents trades, $C$ represents cancellations, and PDM denotes price downward modifications.
[…]
Figure 1: Illustration of oracle construction using passive-aggressive trade pairs within interval $w=[w_{1},w_{2})$. The first-leg execution determines the oracle decision $\chi^{m}(w)$.
Remark.
(Trade-Only Basis for oracle Construction)
Only Trade events are used to construct the oracle decision $\chi^{m}(w)$. Events of type Cancellation or Price Downward Modification are excluded from the filtered set $\mathcal{G}^{\mathcal{T}}(w)$, ensuring that benchmarking reflects executed spread trades alone.
$\tilde{S}_{m}(t)=p_{n}^{b,T}(t_{2})-p_{c}^{a,T}(t_{1})$
, where $T$ represents event-type Trade. The oracle decision is:
[…]
Remark.
(On Quantity Normalization)
We define the oracle label $\chi^{m}(w)$ by sequentially evaluating tick data for calendar-spread trade pairs, i.e., opposite-direction trades across $F_{c}$ and $F_{n}$ within each considered window. For each valid pair, we compute the realized spread using trade-derived prices. If multiple valid pairs occur within $w$, we retain the pair that attains the smallest spread. This yields a hindsight lower bound for evaluation. The leg that trades first in the retained pair determines the oracle decision. The complete filtered set is denoted $\mathcal{G}^{\mathcal{T}}(w)$, and all benchmarking is grounded in this empirically observed structure.
@@ 3.2 Benchmarking Framework for quoting policies
3.3 Tick-based quoting rule.
@@
3.4 Forecasting Execution Risk with Multivariate Hawkes Processes